OpenStudy (anonymous):

if x= y-3 and y= z^2, what is x in terms of z

OpenStudy (mr.math):

Plug \(y=z^2\) into the first equation, what would you get?

OpenStudy (anonymous):

Hint: Substitution

OpenStudy (anonymous):

still confused

OpenStudy (mr.math):

The first equation \(x=y-3\), but y is the same as z^2 by the second equation. So \[x=z^2-3\]

OpenStudy (anonymous):

ughhh still not getting it im trying to solve it myself

OpenStudy (mr.math):

Oh do you know what "x in terms of z" means?

OpenStudy (anonymous):

no

OpenStudy (mr.math):

It means to write x=some expression in z, likw x=z, x=2z+3, x=z^2+1, etc.

OpenStudy (mr.math):

But you still need to satisfy the given equations.

OpenStudy (anonymous):

so is the answer z^-2/3

OpenStudy (mr.math):

The answer is \(x=z^2-3\).

OpenStudy (anonymous):

but that's not one of the options

OpenStudy (mr.math):

Then either your options are all wrong or you didn't write your question right! :D

OpenStudy (mr.math):

Is it \(x=y-3\) or \(x=y^{-3}\)?

OpenStudy (anonymous):

its x=y-3 and y=z^2 and the answers are A z^3/2 B z^5 C z^-6 D z^-2/3

OpenStudy (mr.math):

Then your answer is C, but your question is \(x=y^{-3}\)...

OpenStudy (anonymous):

ok thanks for helping me i appreciate

OpenStudy (mr.math):

Anytime! :)